### v2

# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

Abstract
Federated Learning (FL) enables model training across decentralized devices without sharing raw data, thereby preserving privacy in sensitive domains like healthcare. In this paper, we evaluate Kolmogorov–Arnold Networks (KAN) architectures against traditional MLP across six state-of-the-art FL algorithms on a blood cell classification dataset. Notably, our experiments demonstrate that KAN can effectively replace MLP in federated environments, achieving superior performance with simpler architectures. Furthermore, we analyze the impact of key hyperparameters—grid size and network architecture—on KAN performance under varying degrees of Non-IID data distribution. In addition, our ablation studies reveal that optimizing KAN width while maintaining minimal depth yields the best performance in federated settings. As a result, these findings establish KAN as a promising alternative for privacy-preserving medical imaging applications in distributed healthcare. To the best of our knowledge, this is the first comprehensive benchmark of KAN in FL settings for medical imaging task.
Index Terms: e-health, federated learning, kolmogorov–arnold networks
I Introduction
Recent years have seen major progress in AI for healthcare, showing great potential across many medical uses [1, 2]. At the same time, growing privacy concerns have restricted access to medical data, leading to more interest in methods that protect privacy [3]. Federated Learning (FL) is a new approach that allows training on spread-out datasets without sharing the actual data [4, 5]. This has made FL popular in healthcare since it meets the strict privacy rules for medical data [6, 7]. However, FL performance hinges on the chosen model and its efficiency in distributed settings [8, 9]. Therefore, studying effective model designs for FL remains important for advancing medical applications.
Figure 1: Benchmarking results comparing KAN-1 layer (filled circles) and MLP-3 layer (unfilled circles) performance across SOTA FL methods in an IID setting, plotting Convergence Round versus Top-1 Test Accuracy (%). Circle sizes indicate standard deviation in accuracy.
The Kolmogorov–Arnold Networks (KAN) [10, 11], recently proposed as a potential alternative to Multi-Layer Perceptrons (MLP) [12], can theoretically represent continuous functions without using traditional activation functions like other neural networks. While recent studies have used KAN for various centralized uses [13, 14], research on using them in federated settings is still new, mostly limited to the vanilla FedAvg method or using simple datasets [15, 16, 17]. Current research lacks thorough comparisons of KAN over recent FL methods, which prevents a deeper understanding of their practical use in privacy-sensitive medical settings. Addressing this gap is crucial to fully understand KAN’s potential in complex, spread-out medical data environments. Thus, comprehensive testing of KAN across recent FL methods is essential.
To evaluate KAN’s applicability to FL, we compare their performance against MLP-based models across various cutting-edge FL methods, as shown in Fig. 1. This comparison shows that KAN models, even in their simplest form, perform as well as deeper MLP networks when used with various FL methods in the IID setting. Additionally, we examine KAN architectures from multiple perspectives, investigating the impact of key hyperparameters—specifically grid size and number of hidden layers—when integrated within FL methods in the Non-IID setting. These findings will encourage further exploration of KAN’s wider use in FL, especially for medical imaging tasks requiring AI interpretability.
The main contributions of this paper are as follows:
  • We present the first comprehensive benchmark of KAN in FL settings for medical imaging.
  • We demonstrate that KAN compare favorably with MLP across six state-of-the-art FL methods.
  • We provide empirical evidence of performance patterns that emerge when varying KAN’s key hyperparameters.
  • In addition, we examine whether increasing KAN’s depth through additional hidden layers or expanding its width has greater impact.
The remainder of this paper is organized as follows. In Section II, we introduce federated setting and process. Section III presents our comprehensive benchmark results across various state-of-the-art FL algorithms. Finally, Section IV concludes with remarks.
II Problem and Method
II-A Federated Setting
We consider a FL system comprising $N$ edge devices and a central server as [18]. Each device $n\in\{1,2,\dots,N\}$ holds a private dataset $\mathcal{D}^{n}$, and due to heterogeneous blood-cell type distributions, the overall data union $\mathcal{D}=\bigcup_{n=1}^{N}\mathcal{D}^{n}$ is inherently imbalanced. Consequently, the goal is to learn a global model $\theta_{g}\in\Theta$ without aggregating raw data centrally, thereby preserving privacy [19, 20].
Accordingly, FL seeks to minimize the empirical risk across all devices:
$\min_{\theta_{g}\in\Theta}F(\theta_{g})\;\triangleq\;\frac{1}{N}\sum_{n=1}^{N}f_{n}(\theta_{g})\,$,
where $F$ denotes the global objective aggregated at the server, and each local objective $f_{n}$ on device $n$ is given by:
$f_{n}(\theta_{n})\;\triangleq\;\frac{1}{|\mathcal{D}^{n}|}\sum_{(x_{i},y_{i})\in\mathcal{D}^{n}}\mathcal{L}\bigl{(}\theta_{n};x_{i},y_{i}\bigr{)}\,$,
with $\mathcal{L}$ representing the per-sample loss (e.g., cross-entropy) for predicting label $y_{i}$ from image $x_{i}$.
II-B Federated Learning Process
For clarity and concreteness, we describe the FL framework using the FedAvg method. At each global round $g_{r}\in\{1,2,\dots,G\}$, the central server randomly selects a subset $\mathcal{S}_{g_{r}}\subseteq\{1,2,\dots,N\}$ of edge devices for participation and broadcasts the current global model parameters $\theta_{g}^{(g_{r}-1)}$ to these selected devices.
Each participating device $n\in\mathcal{S}_{g_{r}}$ then performs local training using its private dataset $\mathcal{D}^{n}$ to minimize the local loss function specific to the KAN model. The KAN model computes output $h(x)$ given input $x$ via a Kolmogorov–Arnold representation as follows:
$h(x)=\sum_{j=1}^{J}c_{j}\phi_{j}\left(\sum_{i=1}^{I}w_{ji}\psi_{i}(x_{i})+b_{j}\right)\,$,
where $I$ and $J$ denote the dimensions of input and hidden layers, respectively, $\psi_{i}$ and $\phi_{j}$ represent continuous basis functions of input and hidden layers, and $w_{ji}$, $b_{j}$, and $c_{j}$ are learnable weights and biases and output scalars.
The local training on device $n$ involves minimizing the following local objective through gradient descent [21] for each local epoch $e\in\{1,2,\dots,E\}$:
$\theta_{n}^{(g_{r},e+1)}\leftarrow\theta_{n}^{(g_{r},e)}-\eta\nabla f_{n}(\theta_{n}^{(g_{r},e)})\,$,
where the local loss $f_{n}(\theta_{n})$ for a given batch $\mathcal{B}\subseteq\mathcal{D}^{n}$ is explicitly computed as:
$f_{n}(\theta_{n})=\frac{1}{|\mathcal{B}|}\sum_{(x_{i},y_{i})\in\mathcal{B}}\mathcal{L}\left(h(x_{i};\theta_{n}),y_{i}\right)\,.$
Once local training concludes, each device $n$ sends its updated local parameters $\theta_{n}^{(g_{r},E)}$ back to the central server. The server then aggregates these parameters from participating devices using a parameter-wise averaging:
$\theta_{g}^{(g_{r})}\leftarrow\frac{1}{|\mathcal{S}_{g_{r}}|}\sum_{n\in\mathcal{S}_{g_{r}}}\theta_{n}^{(g_{r},E)}\,$,
where each parameter set $\theta_{n}$ represents set of parameters $\{w_{ji},b_{j},c_{j}\}$ of the KAN model. The aggregated global model $\theta_{g}^{(g_{r})}$ is then broadcasted back to all devices, initiating the next global round. This federated procedure iteratively continues until convergence criteria for the global model $\theta_{g}$ are satisfied. Overall procedure is described in Algorithm 1.
1Initialize global model parameters $\theta_{g}^{(0)}$
2 for global round $g_{r}=1,2,\ldots,G$ do
3   Server randomly selects a subset of devices $\mathcal{S}_{g_{r}}\subseteq\{1,2,\ldots,N\}$
4   Server broadcasts $\theta_{g}^{(g_{r}-1)}$ to all devices in $\mathcal{S}_{g_{r}}$
5
6 for device $n\in\mathcal{S}_{g_{r}}$ do in parallel
7      Initialize local model: $\theta_{n}^{(g_{r},1)}\leftarrow\theta_{g}^{(g_{r}-1)}$
8
9    for local epoch $e=1,2,\ldots,E$ do
10         Sample batch $\mathcal{B}\subseteq\mathcal{D}^{n}$
11         Compute KAN output for each $(x_{i},y_{i})\in\mathcal{B}$:
12       $h(x_{i};\theta_{n}^{(g_{r},e)})=\sum_{j=1}^{J}c_{j}\,\phi_{j}\!\left(\sum_{i=1}^{I}w_{ji}\,\psi_{i}(x_{i})+b_{j}\right)$
13         Compute local loss:
14       $f_{n}(\theta_{n}^{(g_{r},e)})=\frac{1}{|\mathcal{B}|}\!\sum_{(x_{i},y_{i})\in\mathcal{B}}\!\mathcal{L}\!\left(h(x_{i};\theta_{n}^{(g_{r},e)}),y_{i}\right)$
15         Update local model:
16       $\theta_{n}^{(g_{r},e)}\leftarrow\theta_{n}^{(g_{r},e)}-\eta\,\nabla f_{n}(\theta_{n}^{(g_{r},e)})$
17
18      end for
19
20     Device $n$ sends $\theta_{n}^{(g_{r},E)}$ to the server
21
22   end forpar
23
24  Server aggregates parameters:
25 $\theta_{g}^{(g_{r})}\leftarrow\frac{1}{|\mathcal{S}_{g_{r}}|}\sum_{n\in\mathcal{S}_{g_{r}}}\theta_{n}^{(g_{r},E)}$
26
27 end for
28
return $\theta_{g}^{(G)}$
Algorithm 1 FL with Kolmogorov–Arnold Networks
Figure 2: Test accuracy comparison of KAN-1 layer versus MLP architectures (1-3 layers) across six FL algorithms on the blood cell classification dataset with Non-IID distribution ($\alpha=1.0$). Note that shaded regions represent standard deviation across five random seeds.
III Experiment and Results
III-A Experiment Setting
To validate the effectiveness of KAN in federated settings, we conduct experiments on the blood cell classification dataset [22] using both KAN and MLP architectures with varying depths. Note that we employed radial basis function-based FastKAN [23] for our experiments to improve computational efficiency. We examine six state-of-the-art FL algorithms: FedAvg [4], FedDyn [24], FedSAM [25], FedGamma [26], FedSMOO [27], and FedSpeed [28] to comprehensively evaluate performance across different federated optimization techniques. In addition, we simulate a Non-IID data distribution using Dirichlet distribution with $\alpha=1.0$ across $N=100$ edge devices as [29], randomly selecting $10\%$ of devices per global round. We report the mean and standard deviation of test accuracy across five random seeds, and additional hyperparameter settings are available in our open source repository [^1] https://github.com/yjlee22/fkan.
III-B Impact of KAN on FL Performance
In this experiment, we aim to explore the potential of KAN as viable replacements for traditional MLP in federated environments. As shown in Fig. 2, KAN consistently outperforms all MLP configurations, suggesting that KAN could potentially replace MLP-based networks in federated environments. Furthermore, the single-layer KAN not only achieves higher accuracy than single-layer MLP but also surpasses the performance of deeper MLP architectures while exhibiting more stable convergence across different random seeds. This stability is clearly evident from the narrower standard deviation bands in the KAN results compared to MLP. Additionally, this advantage remains consistent across all examined FL optimizers, with certain algorithms showing particularly notable synergy with KAN architecture. Our findings demonstrate that KAN provide superior performance for Non-IID data in federated settings, making them an effective alternative to MLP for medical imaging applications.
Figure 3: Impact of KAN grid size ($g=3,5,10$) on test accuracy across different Non-IID degrees (Dirichlet parameter $\alpha$ from $0.001$ to $1.0$) for six federated learning algorithms. Lower $\alpha$ values indicate higher data heterogeneity across devices. Note that error bars represent standard deviation across five random seeds.
III-C Impact of Non-IID Degree
We investigate how varying the Non-IID degree (Dirichlet parameter $\alpha$) influences the performance of KAN architectures by altering the grid size, as shown in Fig. 3. Results indicate that increasing data heterogeneity (lower $\alpha$) negatively affects test accuracy across all FL methods. In detail, larger grid sizes (e.g., $g=10$) generally yield better performance, suggesting grid size is a critical hyperparameter for mitigating Non-IID data challenges. Notably, the improvement from larger grids becomes more pronounced as data heterogeneity intensifies, highlighting grid size’s importance in highly heterogeneous scenarios. Moreover, specific FL methods, such as FedGamma and FedSpeed, shows notable synergy with larger grid configurations. These findings underscore that appropriately selecting the grid size is essential for optimizing KAN performance in diverse federated learning environments.
TABLE I: KAN variants evaluated in the ablation.
Config Hidden-layer Config Hidden-layer
$d_{1}$ [5] $\omega_{1}$ [5]
$d_{3}$ [5, 5, 5] $\omega_{3}$ [125]
$d_{5}$ [5, 5, 5, 5, 5] $\omega_{5}$ [3125]
III-D Ablation Study
We perform an ablation study using vanilla FedAvg to determine the relative impact of increasing KAN’s depth ($d$) versus expanding its width ($\omega$) as Table. I. As depicted in Fig. 4, results demonstrate that increasing network depth does not guarantee performance improvements, as shallower KAN architectures (1 layer) actually outperform deeper variants (3 and 5 layers) in our experiments. In contrast, width shows significant impact on performance, with the medium configuration (125 parameters) achieving optimal test accuracy compared to both narrow (5 parameters) and wide (3125 parameters) variants. The results indicate the importance of carefully tuning width parameters while keeping depth minimal for enhanced KAN performance in FL scenarios. Note that depth may have a more significant positive impact than observed in our experiments.
Figure 4: Test accuracy comparison of KAN variants with different depths (top) and widths (bottom) across global rounds.
IV Conclusion
In this work, we present the first comprehensive benchmark of KAN in FL for medical image classification. Our results show that KAN can outperform MLP-based network across multiple FL algorithms while maintaining better stability in Non-IID settings. Additionally, we find larger grid configurations provide greater benefits as data heterogeneity increases, offering valuable insights for real-world deployment. Contrary to conventional design principles, our studies reveal that increasing KAN width yields more benefits than increasing depth for medical imaging tasks. Therefore, these findings establish KAN as a promising alternative for privacy-sensitive medical applications, opening new research directions for federated healthcare environments.